• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 16:04:44 on August17,2001

    Entry number 71042

    keyword=shift summary - Aug 17 day

    **Executive summary: problems problems, not a good shift** but the target polarization is high with no beam!

    8:00 AM: taking data as shift starts, A_||, 0 degrees
    schedule for day under discussion

    8:51: ask to take beam away for target polarization measurements; a few
    miutes later MCC calls and tells us (I think) box supply troubles, no beam for a
    while

    9:10 Target pol measurements done, ask for beam back

    ~9:15 MCC looking at injector for source of current instability

    ~10:10 check with MCC, still working on it, probably another hour

    ~10:50 beam back for checks, with raster off, and BPM positions reading
    x = -0.055, y=-1.605, beam appears centered on the BeO cross hairs.

    suddenly noticed beam at y = -4.5, not sure how it got there, MCC says they
    just noticed it and are resteering.

    11:10 AES looking RF problem, no estimate

    1:10 high temperature in magnet: air conditioner failed, new one being
    rolled over, maybe beam at 2PM

    2:25 MCC says they are ready to deliver beam. Ask for 5 uA on BeO.

    2:35 raster looks okay, doing target polarization before taking data

    2:55 just done with target polarization and rotating field, ready for beam.
    Call MCC and find they just seconds before lost another magnet due to the heat.

    3:40 MCC calls, asking for what beam we want

    4:00 get ~ 2 minutes of beam, coda daq problem to fix during this...